The expected pretax return on three stocks is divided between dividends and capital gains in the following way:

Stock Expected Dividend Expected Capital Gain

A $0 $10

B $5 $5

C $10 $0

a) If each stock is priced at $120, what are the expected net percentage returns on each stock (i) a pension fund that does not pay taxes, (ii) a corporation paying tax at 35% (the effective tax rate on dividends received by corporations at 10.5%), and (iii) an individual with an effective tax rate on 15% on dividends and 10% on capital gains? (Do not round immediate calculations. Enter your answers as a percent rounded to 2 decimal places).

Stock Pension Investor Corporation Individual

A _____% _________% __________%

B ______% _________% __________%

C _______% _________% __________%

b) Suppose the investors pay 50% tax on dividends and 20% tax on capital gains. If stocks were priced to yield an after-tax return on 8%, what would A, B, and C each sell for? Assume that the expected dividend is a level perpetuity. (Do not round immediate calcuations. Round your answer to 2 decimal places).

Stock Price

A ________

B ________

C __________
Answer #1

Answer (a):

Individual Pension Investor Corporation STOCK 8.33% 5.42% 7.50% 8.33% 6.44% 7.29% 8.33% 7.46% 7.08%

Working:

(i) Pension Fund Expected Dividend (D) Expected capital gain (CG) $10 Tax on dividend Tax on Capital Gain Total after tax % N

Answer (b):

After tax return required = 8%

Capital gain tax =20%

Tax on dividend = 50%

Stock A

After tax dividend = 0 * (1- 50%) = $0

After tax capital gain = $10 * (1- 20%) = $8

Selling Price = ($0 + $8) / 8% = $100

Stock B

After tax dividend = 5 * (1 - 50%) = $2.50

After tax capital gain = $5 * (1- 20%) = $4

Selling Price = ($2.50 + $4) / 8% = $81.25

Stock C

After tax dividend = 10 * (1 - 50%) = $5

After tax capital gain = $0 * (1- 20%) = $0

Selling Price = ($5 + $0) / 8% = $62.50

Hence:

A Selling price = $100

B Selling price = $81.25

C Selling price = $62.50